Although appealing to more refined tastes, art as a collectible has not always performed so profitably. Assume that in 2015, an auction house sold a statute at auction for a price of $10,584,500. Unfortunately for the previous owner, he had purchased it in 2008 at a price of $12,624,500. What was his annual rate of return on this sculpture?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

10,584,500=12,624,500*(1+r/100)^7

(10,584,500/12,624,500)^(1/7)=1+r/100

1+r/100=0.975135946

r=0.975135946-1

=(2.49%)(Approx)(Negative).